According to indexmundi.com China produces the largest amount of ammonia with 45,520 in thousand metric ton. The profitability of the ammonia production is around 90% of the total ammonia production cost which is about Php 10,000 per ton.
Refer to the balanced chemical equation: N2(g) + 3H2(g) ➡️ 2NH3(g)

How much they gain from this amount of ammonia produced?
1
Expert's answer
2020-11-26T04:53:47-0500

45520 metric tons= 50177tons

1 ton costs Php 10,000

50177 tons = 10000×50177 = Php 501,770,000

Profit = 90/100 × 501,770,000 = Php 451,593,000